OpenStudy (anonymous):

Please help ASAP!!! Write the following numbers, in order, from least to greatest. Write the numbers with a comma between each. -2, 1.5, -1, -0.89, 1.75

OpenStudy (anonymous):

-2, -1, -0.89, 1.5, 1.75

OpenStudy (confluxepic):

You put the negative numbers first and then the positive numbers in order.

OpenStudy (confluxepic):

The median is the middle number which is -.89

OpenStudy (anonymous):

Thank you could you please help me with another?

OpenStudy (confluxepic):

Sure.

OpenStudy (confluxepic):

I will try.

OpenStudy (anonymous):

1/2 <,>,= to 3/7

OpenStudy (confluxepic):

First you find the values.

OpenStudy (confluxepic):

1/2=.5

OpenStudy (confluxepic):

3/7=.42

OpenStudy (confluxepic):

1/2 is greater than .42 so it would look like 1/2>3/7

OpenStudy (anonymous):

Thank you so much! Could you help me with one more please?

OpenStudy (confluxepic):

Sure.

OpenStudy (anonymous):

 Least to greatest.. I know 0 would be first...